OpenStudy (anonymous):

can any1 help me out....A random number generator draws at random with replacement from the digits 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. In 5000 draws, the chance that the digit 0 appears fewer than 495 times is closest to a)25% b)30% c) 35% d) 40% e) 45% f) 50%

OpenStudy (zarkon):

This is a binomial distribution problem

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Normal approximation to binomial p = 0.1 of getting a zero mean E(x) = np = 0.1*5000 = 500 SD = sqrt(np(1-p)) = sqrt(500*0.9) = 21.21 for normal approximation z = (494.5-500)/21.21 = -0.26 P(x < 495) = P(X < 494.5) = 39.7%

OpenStudy (anonymous):

got it...:)...thanks dear..